OpenStudy (anonymous):

HELP ME PLEASE I HAVE A FINAL TOMORROW!!!!! For the following logarithmic function, determine the domain, vertical asymptote, and the x-intercept f(x)=ln(x+1) PLEASE, I NEED YOUR HELP, PLEASE

ganeshie8 (ganeshie8):

which part are you stuck at ?

ganeshie8 (ganeshie8):

`domain` is the set of all values for which the function is defined. logarithmic function is defined only for positive numbers, so simply solve below for domain : \[\large x+1\gt 0\]

OpenStudy (anonymous):

x>-1 therefore -1,infinity

OpenStudy (anonymous):

ok there goes the domain and the VA is x=-1 cause of just solving it

OpenStudy (anonymous):

ok so I got domain and VA

OpenStudy (anonymous):

NOW, how do you do Intercept again???

ganeshie8 (ganeshie8):

Perfect! Domain : (-1, \(\infty\)) VA : x = -1

OpenStudy (anonymous):

thank-you :)

OpenStudy (anonymous):

ok so anyways, intercept???

ganeshie8 (ganeshie8):

x intercept is the x value that makes the function equal to 0

ganeshie8 (ganeshie8):

solve below for x intercept : \[\ln (x+1) = 0\]

OpenStudy (anonymous):

isn't it -1???

ganeshie8 (ganeshie8):

plugin x = -1 into the function, does it spit out 0 ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

careful

ganeshie8 (ganeshie8):

\[\ln (x+1) = 0\] is equivalent to \[x+1 = e^0\] which is same as \[x+1 = 1\] \[x = ?\]

OpenStudy (anonymous):

0

ganeshie8 (ganeshie8):

Yep!

ganeshie8 (ganeshie8):

you may get more convinced by looking at the graph https://www.desmos.com/calculator/bs9qbr0fgo

OpenStudy (anonymous):

Ok um it said that the Intercept was (0,0) not just 0, so may you explain that please???????

OpenStudy (anonymous):

the graph just answered my question, didn't it?

ganeshie8 (ganeshie8):

yes, x = 0 is a number (0, 0) is a point

ganeshie8 (ganeshie8):

your textbook is saying the x intercecpt point is (0, 0) which should make sense, eh ?

OpenStudy (anonymous):

ohhhhh, I see, thank-you very much and the answer sheet my teacher gave me said that btw

OpenStudy (anonymous):

ok and if I may I just just ask how to find the range of this equation:

OpenStudy (anonymous):